OpenStudy (anonymous):

what is the area of the circle that has the equation below: x^2+y^2+2x-12y+12=0 5pi 12pi 25pi 144pi

OpenStudy (anonymous):

you'll need to put this general equation of a circle in center-radius form by completing the square... then you'll have the radius in which case you can calculate the area

OpenStudy (anonymous):

group the expressions with x and the expressions with y together....

OpenStudy (anonymous):

x2+2x+y2-12y=-12 x^2+2x+1-1+y2-12y+36-36=-12 (x+1)^2+(y-6)^2=-12+36+1 =25 5^2 radius =5 area =2pie 5

OpenStudy (anonymous):

area =pie 25 make correction